OpenStudy (anonymous):

If a ball is thrown straight up into the air with an initial velocity of 80 ft/s, its height in feet after t second is given by y=80t−16t^2. a) Find the average velocity for the time period begging when t=2 and lasting (i) 0.1 seconds...... i.e., over the time interval [2,2.1] (ii) 0.01 seconds (iii) 0.001 seconds b)Estimate the instantaneous velocity of the ball when t=2.

OpenStudy (anonymous):

i got 14.4 for i

OpenStudy (anonymous):

radar i think i got it all except b

OpenStudy (anonymous):

actually (iii), i need help with that

OpenStudy (radar):

80-64=14 fps.......wouldn't that be the answer the initial velocity minus the reduced velocity (due to gravity)

OpenStudy (radar):

using the derivative of dh/dt d' of 80t-16t^2 = 80-32t

OpenStudy (radar):

plug in t=2.

OpenStudy (radar):

I think part a is the most difficult, good luck with this

OpenStudy (anonymous):

got it thanks

OpenStudy (anonymous):

radar do you about limits?

OpenStudy (anonymous):

as x is approaching a number, that type of limit?

OpenStudy (radar):

not to much, I do agree with you on the 14.4 for a i for the limit please state the function of x and see if I can figure it out

OpenStudy (anonymous):

looking at this graph, im supposed to determine a bunch of limits first x as it approaches 1 from left then as x approaches 1 from right as x approaches 1 f(1) as x approaches 2 from left as x approaches 2 from right as x approaches 2 f(2)

OpenStudy (radar):

Sorry, but not much help here the function looks to be discontinous at a couple of points. Better post that and get a lot exposure for it lol

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

what about this Given that f(x)=x^2−7x and g(x)=x−5, calculate f of g g of f f of f g of g

OpenStudy (radar):

Here is my wag on f(1) 0 as x approaches 1 from the left 0 as x approaches 1 from the right 0 as x approaches 1 f(2) 2 as x approaches 2 from the left 1 as x approaches 2 from the right indeterminate at 2

OpenStudy (radar):

f of g = (x-5)^2-7(x-5) x^2-10x+25-7x+35 x^2-17x+60

OpenStudy (anonymous):

wow you were right with all. thanks very much